OpenStudy (anonymous):

rationalize the denominator -11x^5 / sqrt(-6x+9)

OpenStudy (anonymous):

\[-11x ^{5}\div \sqrt{-6x+9}\]

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{-11x^5}{ \sqrt{-6x+9}} }\)

OpenStudy (solomonzelman):

multiply top and bottom times √{-6x+9}

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{-11x^5\color{red}{\times\sqrt{-6x+9}}}{ \sqrt{-6x+9}\color{red}{\times\sqrt{-6x+9}}} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{-11^5\sqrt{-6x+9}}{-6x+9} }\)

OpenStudy (anonymous):

and it can't be simplified further? that's what i got!

OpenStudy (solomonzelman):

And btw, domain ≤ 2/3

OpenStudy (solomonzelman):

No it can't be simplified further.

OpenStudy (anonymous):

thank you!

OpenStudy (solomonzelman):

Well, you could have said, instead of rationalizing, \(\normalsize\color{blue}{ -11x^5(-6x+9)^{-1/2}}\) but this is not rationalized. You aren't looking for that.

OpenStudy (solomonzelman):

You welcome !